OpenStudy (anonymous):

Find equations of (a) th tangent plane and (b) the normal line to the given surface at the specific point. A) 2(x-2)^2+(y-1)^2+(z-3)^2=10, P=(3,3,5)

hero (hero):

I can get the answer, but it may take a bit.

OpenStudy (anonymous):

yeh great, take your time...

hero (hero):

I'll be back

hero (hero):

Somebody may give the solution before I get back. If that happens, that's cool too.

OpenStudy (anonymous):

ok I hope so, I really want to understand this

OpenStudy (anonymous):

I think I got it! This is the solution for whoever would like to learn more about this topic: First, we find the gradient of the Surface: which is, del S = <4(x-2), 2(y-1), 2(z-3)> then at the point P=(3,3,5), we get the gradient is equal to: del S at (3,3,5)= <4,4,4> We now can find the quation of the plane: 4(x-3)+4(y-3)+4(z-5)=0 Simplify: 4x-12+4y-12+4z-20=0 4x+4y+4z=44 x+y+z=11 <-- that is the equation of the plane and the normal line is: <1,1,1> equation to the normal line in the symmetric form: X-3=y-3=z-5 Thanks Hero anyways, if you come back and you see this.